gelphielvr:

(algebra 1) inequalities. question in the replies

gelphielvr:

1 attachment
SoundwaveKun:

Do you know how to start this?

SoundwaveKun:

I guess not, I'll help you! 1. Distribute the 5 on the left side 2. Move all terms involving x to one side and constant terms to the other side. Start by adding 5x5x5x to both sides 3. Next, subtract 6 from both sides 4. Divide both sides by 3 Then you should have your answer! ^_^

gelphielvr:

@soundwavekun wrote:
I guess not, I'll help you! 1. Distribute the 5 on the left side 2. Move all terms involving x to one side and constant terms to the other side. Start by adding 5x5x5x to both sides 3. Next, subtract 6 from both sides 4. Divide both sides by 3 Then you should have your answer! ^_^
sorry I was offline. do you know if in the end I have to flip the signs?

ThanosQC:

1) 5(3-x) < -2x + 6 -Distribute the "5" 5*3 = 15, -x*5 = -5x 2) 15-5x < -2x + 6 -Isolate for x 15-6-5x+5x < -2x+5x+6-6 9 < 3x 3) -Divide and solve for x 9/3 < 3x/3 3 < x, so when properly formatted, x > 3

gelphielvr:

@thanosqc wrote:
1) 5(3-x) < -2x + 6 -Distribute the "5" 5*3 = 15, -x*5 = -5x 2) 15-5x < -2x + 6 -Isolate for x 15-6-5x+5x < -2x+5x+6-6 9 < 3x 3) -Divide and solve for x 9/3 < 3x/3 3 < x, so when properly formatted, x > 3
thank you :)